# Funding AI for Good: A Call for Meaningful Engagement

### v2

@@ 3.3. Positionality
The authors’ collective prior research and practice in public sector AI and AI for social good have led us to believe that funding agendas are important to study. We started with the basis that power imbalances exist within AI4SG partnerships and sought to examine power from the very beginning of AI4SG projects: when they were conceptualized for funding. While the authors are from diverse ethnic and cultural backgrounds, all are researchers in academia living in the United States and conduct research in English.
@@ 4. Findings
In what follows, we report on themes in different components of funding documents and highlight rhetoric leaning towards a more techno-centric or balanced approach to AI4SG. We begin with the Background (Section 4.1), where we discuss how funding documents framed “social good”, “AI”, and how AI is envisioned to help address social good problems (RQ1). In the Expected Outcomes and Evaluations (Section 4.2), we discuss funding documents characterized outcomes, often communicated through evaluation criteria (RQ2). In the Eligibility (Section 4.3) and Support from Funders (Section 4.4) sections, we shed light on how funding documents described “AI4SG” can be achieved and by whom (RQ3). Table LABEL:tab:funding_doc_mr in Appendix A provides a detailed overview of our dataset, including funding amount, duration, targeted regions, funder types, and target regions.
@@ 4.1. Background: Motivations and Beliefs Underlying AI4SG Funding Documents
Most AI4SG funding documents in our corpus focused on addressing large-scale, complex social issues, including public health (6, e.g. [C17, C19]), development (6, e.g. [A7, A8]), climate change (3, e.g. [C11, C12]), and mental health (2, [C8, C9]). These social problems have been categorized as “wicked problems”—complex, large-scale, and often ill-defined problems created by multiple, interdependent, evolving issues that any single intervention cannot fully address (Gurumurthy and Chami, 2019; Lönngren and van Poeck, 2021). However, some funding documents depicted AI-based interventions as a solution to these large-scale social problems. They emphasized AI’s transformative benefits deterministically while glossing over or omitting the specific contexts and stakeholders that AI could be beneficial for (Sections 4.1.1). Some funding documents, on the other hand, were grounded in a specific context within a large-scale social issue, sometimes providing a detailed literature review. When describing AI’s role in addressing the social problems, some funding documents acknowledged both the potential benefits and harms of AI in these specific contexts (Section 4.1.2).
@@ 4.1.1. More techno-centric
Emphasizing AI’s transformative potential with limited contextual grounding. Funding documents in our corpus rarely provided an explicit definition of what “AI” is 222with the exception of C3 and C13 (both in the Frequently Asked Questions sections). Yet several funding documents emphasized AI’s role as a potential solution to complex (and often vaguely defined) social issues, without instantiating exactly how AI could be beneficial and for what contexts. For example, UNICEF Innovation Fund Call for Data Science & A.I explicitly looked for “data science, machine learning, artificial intelligence or similar technology solutions” that would broadly “benefit humanity” (C2). The call provided example questions to elaborate on what benefiting humanity with AI means, though the application area remains broad and vague while pre-determining the use of AI as a solution: “Are you using new sources of data such as satellite imagery or social media and applying data science or artificial intelligence techniques to understand the physical world that we live in?” Under this question, the call then further listed out examples of how AI might be beneficial, such as “[a]pplying deep learning to analyze satellite imagery and automatically map infrastructure such as schools, health centers, roads or cell towers.” While this example provided a more specific example of how AI could be used for a specific task (using satellite imagery to automatically map infrastructure), the call did not make clear why automatically mapping infrastructure like schools might be beneficial, and for whom it would be beneficial.
@@
Promoting urgent dissemination of AI to resource-constrained contexts. Several AI4SG funding programs were designed to diffuse the benefits of AI technologies to developing countries and the nonprofit sectors from Western countries and the private technology sector. This diffusion of AI was seen to be “urgent,” so that under-resourced communities in developing countries or the nonprofit sectors could keep up with the rapid development of AI. Some funding programs from international development agencies in Western countries, such as the AI for Development Africa program funded by IDRC (Canada) and SIDA (Sweden), emphasized the importance of ensuring “developing countries are not left behind in the AI revolution” (A9). Some funding programs offered by private tech companies and philanthropic foundations similarly extend this narrative to the nonprofit sectors: “While private-sector businesses have been building and deploying artificial intelligence for years, most organizations in the nonprofit, civic, and public sectors have yet to robustly apply these techniques towards the complex challenges they address” (A6). As Patrick J. McGovern Foundation’s grant announcement (A1) stated, enabling nonprofits to use AI and data “have never been more urgent.” In these funding documents, the rhetoric of AI4SG led with the assumption that AI’s benefits could be directly transferred to non-Western countries and the nonprofit sectors.
@@ 4.2.1. More techno-centric
“Demonstrating” the benefits of AI with research outputs. Other funding programs in our dataset aimed at advancing use-inspired AI research (without a requirement for AI deployment), including Amazon’s Data for Social Sustainability grant (C6) and Microsoft’s AI for Earth program (A12 and A13), both of which came from the research arms of private tech companies. Most funding documents expected the dissemination of research outputs and contributions to the research community, often in the form of “publications, presentations, code and data releases, blogs/social media posts, and other speaking engagements” (C6), while placing limited to no emphasis on disseminating research outputs with contributors or partners in the community. Moreover, some funding programs sought to demonstrate rather than evaluate the uses of new AI technologies. For instance, the press announcement for the AI for Earth program (A12) described that they intended to “demonstrate how AI can deliver results more rapidly, accurately, and efficiently” and “encourage others to innovate based on the power and potential of AI.” This statement was grounded in the assumption that AI would be deterministically beneficial.
@@ 4.2.2. More balanced
Evaluating AI’s impact and encouraging flexible expectations of novelty in research. Some funding calls aimed at advancing AI research explicitly supported assessments on the direction of AI’s potential impact (rather than just the scale of its benefits). For example, the Wellcome Mental Health Data Prize: UK and South Africa (C8) encouraged applicants to consider the risks posed by data biases in AI technologies and requires applicants to address data quality, bias mitigation, and ethics in their proposals. Applicants must consider “risks or issues associated with the proposed timelines and plans for mitigation,” with this criterion accounting for 10% of applicant evaluation.
Some funding calls also expected research novelty more flexibly to ensure the exploration of AI-based interventions was suitable for the specific contexts. For example, the Wellcome Mental Health Data Prizes (C8 and C9) clarified that “it is not expected that the team must innovate in their methodology if this is not appropriate for their research.” Similarly, the Climate Change AI Innovation Grants (C13) specified that “it is perfectly acceptable to use a previously published ML algorithm or architecture,” and grounded the exploration of AI on a solid contextual understanding: “Projects should employ or address AI/ML in a way that is well-motivated and well-scoped for the problem setting, which may or may not include the development of novel techniques.”
@@ 4.3. Eligibility Criteria and Team Composition
Funding programs in our corpus commonly viewed interdisciplinary collaborations as valuable for achieving the aims of AI4SG. However, some funding calls had requirements that were biased towards (and could only be realistically met by) teams with existing AI capacities (Section 4.3.1). Funding calls also differed in how they scaffolded involvement of different stakeholders, as well as who was seen as a “domain expert” (e.g., those with scholarly knowledge versus lived experiences). Overall, we observed that many funding documents that encouraged involvement of community stakeholders centered around consultative modes of participation and during later AI project stages (i.e., implementation phase) (c.f. (Delgado et al., 2023)) (Section 4.3.1). In Section 4.3.2, we spotlight a few funding documents that went beyond the consultative mode and encourage community ownership.
@@ 4.3.1. More techno-centric
Community engagement limited to consultation in later stages of projects. Across many funding documents in our corpus, we observed that expectations for community engagement—in cases where they existed at all—seldom surpassed the consultation mode, where impacted communities and intended end beneficiaries could provide input, but they were not expected to shape the projects’ overall scope and purpose or play a central role across the project life cycle. For example, some funding calls only encouraged community consultation in the implementation stage, rather than the early stages of problem identification and solution development. For example, the Google AI Impact Challenge (C3) specifically asked applicants to “identif[y] the right partners and domain experts needed for implementation.” The funding call explained that involving these stakeholders during implementation could help “deploy the AI model [to support] real-world impact,” where expected community engagement primarily served the purpose of AI deployment.
@@ 4.4. Funders’ Support and Connections
Funding calls in our dataset supported projects ranging from three months to five years (with a median of one and a half years). The longer-term programs were typically funded by governmental agencies (e.g., C17 and C18 by NIH, and C19 by NSF). Overall, we observed a lack of provisions for post-deployment funding support. Many funding programs readily provided non-monetary support that was biased towards technical aspects of the projects (Section 4.4.1). Some other funding programs provided broader forms of support beyond the technical, sometimes through partnerships with regional organizations (Section 4.4.2). However, across the funding documents in our corpus, there was a notable lack of guidance and training for applicants to develop skills in community and stakeholder engagement.
@@ 4.4.2. More balanced
Partnering with regional organizations where the projects are based. A few funding programs involved regional partners in the design of the funding calls (C8-C12). A notable example was the design process of Wellcome Trust’s Mental Health Data Prizes (C8 and C9). After the first round of the data prize in the UK and South Africa (C8), the funder reflected that “it did not attract as many applications as expected from South Africa” (Wellcome Trust, 2023b). After conducting feedback sections (Wellcome Trust, 2023a), the funder worked with a regional partner, the African Population and Health Research Center, to develop the funding call in Africa (C9). The regional partner played a “primary role” in “prize management” (e.g., developing thematic areas and leading the application selection process), “capacity building” (e.g., providing training in data and mental health research), “matchmaking” (e.g., reach communities and potential participants in Africa), and “landscape analysis of existing mental health research” (e.g., literature review on existing data sources) (C9). The partnership with the regional partner resulted in several key adjustments: providing a literature review on mental health issues in the African context, extending the overall timeline for project development, and indicating that a “lived experience expert” would be a member of the application selection panel.
@@ 5. Discussion
Our findings extend prior literature on document analysis of AI imaginaries (Zhong et al., 2025), by showing that some AI4SG funding documents similarly reproduce many of the same techno-centric assumptions found in ethics toolkits (Wong et al., 2023b), national AI policies (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022), and military funding solicitations (Widder et al., 2024), but with distinct implications for social-impact work. Like public texts that portray AI as an “autonomous and inevitable force” (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022; Widder et al., 2024), some AI4SG funding calls in our corpus similarly positioned AI as a transformative solution to large-scale social problems (with limited contextual grounding) and promoted its urgent dissemination to resource-constrained social contexts (like the nonprofit sector and international development) (Section 4.1.1). Some funding documents also mirrored AI ethics toolkits that “frame the work of AI ethics as technical work for individual technical practitioners” (Wong et al., 2023b), providing little to no guidance on how to involve community stakeholders despite encouraging interdisciplinary collaborations. Yet AI4SG funding documents differ from these domains in their explicit aim to advance social good, making the dissonances between expected outcomes and some techno-centric rhetoric particularly stark. We discuss these dissonances in Section 5.1.
@@
At the same time, a subset of AI4SG documents in our dataset adopted more balanced framings, such as situating AI’s potential within contextual understanding of scoped social problems (Section 4.1.2) and requiring domain expertise or contextual knowledge in core project teams (Section 4.3.2). These examples reveal the presence of funding rhetoric that does encourage community-grounded approaches, a departure from dominant AI imaginaries. While there is complexity in how what is written in funding documents influences what actually gets funded, funding rhetoric influences downstream applicants’ behaviors (Peng et al., 2024) and research activities (Widder et al., 2024; Smith et al., 2023). This makes intervening in the design process of funding documents crucial for influencing downstream project approaches before projects are even selected. Building on the more balanced approaches and existing gaps identified in our dataset, we propose opportunities for designing future funding calls for funders, including supporting relationship building, maintenance post-project deployment, and community engagement training for project teams (Section 5.2). Finally, we outline an agenda for meaningful community engagement for funders, and specify the roles that the HCI researchers and practitioners can and should play in helping shape AI4SG funding agendas (Section 5.3)
@@
5.1. Revealing Dissonances Between Expected Outcomes and More Techno-centric Approaches in AI4SG Funding Documents
@@
Many AI4SG funding documents in our dataset expected feasible and sustainable deployment of AI technologies or novel research advancements (Section 4.2). However, more techno-centric approaches promoted in some funding documents may influence downstream approaches taken by AI4SG project teams, impacting AI4SG’s goal of achieving positive and long-lasting change in real-world contexts (Perera, 2024; Tudor et al., 2024). We highlight three prominent aspects below, in conversation with HCI literature on “tech for good” projects.
@@
First of all, funding documents motivating AI’s use for addressing large-scale social problems without instantiating concrete contextual benefits (Section 4.1.1) could lead to premature emphasis on scalability, risking putting such a strong emphasis on creating something that works everywhere that it ends up not working particularly well anywhere. As Hanna and Park note, “scale thinking” necessitates homogenizing users into predefined categories, disregarding the diversity of local contexts (Hanna and Park, 2020). The urgency to scale could incentivize rushed implementations that are poorly suited to diverse contexts (Heeks, 2003, 2002; Dada, 2006). As Burrell and Toyama (2009) put it, “in an effort to solve problems quickly on large scale, a concern for local values, preferences, and opinions may become secondary.” We do not argue that scalable and generalizable approaches should never be pursued, but we suggest that they should be pursued inductively. That is, AI4SG initiatives should focus early efforts on solving specific problems of specific communities effectively and sustainably, and then look for commonalities that could inform the development of generalizable technical capabilities and infrastructures.
@@
Secondly, funding rhetoric that led with assumptions of AI’s deterministic benefits without acknowledging its risks (Section 4.1.1) could reinforce the wider hype and “enchanted determinism” surrounding AI (Campolo and Crawford, 2020). This narrative fosters unrealistic expectations (Lin et al., 2024; Taylor et al., 2013), promotes techno-solutionism, and diverts attention from underlying social issues (Walsham, 2017; Karusala and Kumar, 2017). Furthermore, when funding documents framed AI as a force for good in the Global South or nonprofit sectors deterministically (Section 4.1.1), it risks promoting top-down approaches that cast these groups as lacking and in need of catching up (Erete et al., 2023; Valencia, 1997; Dearden and Tucker, 2016). This framing ignores the unique needs (and assets) of public sectors and the Global South (Hanna and Park, 2020). While corporate donations of computing resources and support (Section 4.4.1) can expand access, they may also reinforce Big Tech’s hegemonic narrative (Collins, 2000; Lukacz, 2024; Nost and Colven, 2022; Seger et al., 2023), addressing surface-level gaps rather than deeper structural issues (Ho et al., 2009b; Burrell and Toyama, 2009; Dearden and Tucker, 2016). Notably, all the funders in our dataset are headquartered in the Global North, even though many funding programs were aimed at applicants or projects located in the Global South. This geographic asymmetry reveals an uneven distribution of hegemonic power (Collins, 2000), as Global North institutions retain disproportionate authority to set funding agendas. These initiatives risk reproducing patterns of saviorism, positioning the Global South as a testing ground for foreign-owned systems while extracting valuable data (Damilare Dosunmu, 2025). These projects also overlook the uneven geography of the AI supply chain; the Global South is already part of AI, but primarily through the underpaid labor that makes AI systems work (Posada, 2022).
@@
5.2. Reframing AI4SG Funding: Recommendations for Funding Call Designs
Building on the more balanced approaches and existing gaps we identified in our dataset, we offer the following recommendations for funding call designs, putting them in conversation with existing literature on “tech for good” projects. We structure these recommendations around the four key components of funding calls, to serve as a template to inform the design of future funding calls for funders, as well as future HCI research directions in Community-Collaborative Approaches (CCA).

### v3

@@ 3.3. Positionality
The authors’ collective prior research and practice in public-sector AI and AI for social good have led us to believe that funding agendas are worth studying. We started with the basis that power imbalances exist within AI4SG partnerships and sought to examine power from the very beginning of AI4SG projects: when they were conceptualized for funding. While the authors are from diverse ethnic and cultural backgrounds, all are researchers in academia living in the United States and conduct research in English.
@@ 4. Findings
In what follows, we report on themes in different components of funding documents and highlight rhetoric leaning towards a more techno-centric or balanced approach to AI4SG. We begin with the Background (Section 4.1), where we discuss how funding documents framed “social good”, “AI”, and how AI is envisioned to help address social good problems (RQ1). In the Expected Outcomes and Evaluations (Section 4.2), we discuss funding documents characterized outcomes, often communicated through evaluation criteria (RQ2). In the Eligibility (Section 4.3) and Support from Funders (Section 4.4) sections, we shed light on how funding documents described “AI4SG” can be achieved and by whom (RQ3).
@@ 4.1. Background: Motivations and Beliefs Underlying AI4SG Funding Documents
Most AI4SG funding documents in our corpus focused on addressing large-scale, complex social issues, including public health (6, e.g. [C17, C19]), development (6, e.g. [A7, A8]), climate change (3, e.g. [C11, C12]), and mental health (2, [C8, C9]). These social problems have been categorized as “wicked problems”—complex, large-scale, and often ill-defined problems created by multiple, interdependent, evolving issues that any single intervention cannot fully address (Gurumurthy and Chami, 2019; Lönngren and van Poeck, 2021). Funding documents also rarely provided an explicit definition of what “AI” is.444with the exception of C3 and C13 (both in the Frequently Asked Questions sections) However, some funding documents readily depicted AI-based interventions as a solution to these large-scale social problems, emphasizing AI’s transformative benefits deterministically while glossing over or omitting the specific contexts and stakeholders that AI could be beneficial for (Sections 4.1.1). Some funding documents, on the other hand, were grounded in a specific context within a large-scale social issue, sometimes providing a detailed literature review. When describing AI’s role in addressing the social problems, some funding documents acknowledged both the potential benefits and harms of AI in these specific contexts (Section 4.1.2).
@@
Themes along a spectrum of more techno-centric to more balanced approaches related to motivations and backgrounds of funding documents in our corpus. More techno-centric: Emphasizing AI’s transformative potential with limited contextual grounding; Promoting urgent dissemination of AI to resource-constrained contexts. More balanced: Situating AI’s potential within contextual understanding of scoped social problems; Framing AI as one tool among many and acknowledging its potential benefits as well as risks.
@@ 4.1.1. More techno-centric
Emphasizing AI’s transformative potential with limited contextual grounding. Several funding documents emphasized AI’s role as a potential solution to complex (and often vaguely defined) social issues, without instantiating exactly what AI is, how AI could be beneficial, and for whom and what purposes. For example, UNICEF Innovation Fund Call for Data Science & A.I. explicitly looked for “data science, machine learning, artificial intelligence or similar technology solutions” that would broadly “benefit humanity” (C2). The call provided example questions to elaborate on what benefiting humanity with AI means, though the application area remains broad and vague while pre-determining the use of AI as a solution: “Are you using new sources of data such as satellite imagery or social media and applying data science or artificial intelligence techniques to understand the physical world that we live in?” Under this question, the call then further listed out examples of how AI might be beneficial, such as “[a]pplying deep learning to analyze satellite imagery and automatically map infrastructure such as schools, health centers, roads or cell towers.” While this example provided a more specific example of how AI could be used for a specific task (using satellite imagery to automatically map infrastructure), the call did not make clear why automatically mapping infrastructure like schools might be beneficial, and for whom it would be beneficial.
@@
Promoting urgent dissemination of AI to resource constrained contexts. Several AI4SG funding programs were designed to diffuse the benefits of AI technologies to developing countries from Western countries, and to the nonprofit sector from the private technology sector. This diffusion of AI was seen as “urgent,” so that under-resourced communities in developing countries or the nonprofit sector could keep up with the rapid development of AI. Some funding programs from international development agencies in Western countries, such as the AI for Development Africa program funded by IDRC (Canada) and SIDA (Sweden), emphasized the importance of ensuring “developing countries are not left behind in the AI revolution” (A9). Some funding programs offered by private tech companies and philanthropic foundations similarly extend this narrative to the nonprofit sectors: “While private-sector businesses have been building and deploying artificial intelligence for years, most organizations in the nonprofit, civic, and public sectors have yet to robustly apply these techniques towards the complex challenges they address” (A6). As Patrick J. McGovern Foundation’s grant announcement (A1) stated, enabling nonprofits to use AI and data “have never been more urgent.” In these funding documents, the rhetoric of AI4SG led with the assumption that AI’s benefits could be directly transferred to non-Western countries and the nonprofit sectors.
@@ 4.2. Expected Outcomes and Evaluations
Themes along a spectrum of more techno-centric to more balanced approaches related to expected outcomes and evaluations of funding documents in our corpus. More techno-centric: Grounding evaluation in technical dimensions over community outcomes; “Demonstrating” the benefits of AI with research outputs. More balanced: Grounding evaluation in benefits to the local community beyond the technical aspects; Evaluating AI’s impact and encouraging flexible expectations of novelty in research.
@@ 4.2.1. More techno-centric
“Demonstrating” the benefits of AI with research outputs. Other funding programs in our dataset aimed at advancing use-inspired AI research (without a requirement for real-world AI deployment), including Amazon’s Data for Social Sustainability grant (C6) and Microsoft’s AI for Earth program (A12 and A13), both of which came from the research arms of private tech companies. Most funding documents expected the dissemination of research outputs and contributions to the research community, often in the form of “publications, presentations, code and data releases, blogs/social media posts, and other speaking engagements” (C6), while placing limited to no emphasis on disseminating research outputs with contributors or partners in the community. Moreover, some funding programs sought to demonstrate rather than evaluate the uses of new AI technologies. For instance, the press announcement for the AI for Earth program (A12) described that they intended to “demonstrate how AI can deliver results more rapidly, accurately, and efficiently” and “encourage others to innovate based on the power and potential of AI.” This statement was grounded in the assumption that AI would be deterministically beneficial.
@@ 4.2.2. More balanced
Evaluating AI’s impact and encouraging flexible expectations of novelty in research. Some funding calls aimed at advancing AI research explicitly supported assessments on the direction of AI’s potential impact (rather than just the scale of its benefits). For example, the Wellcome Mental Health Data Prize: UK and South Africa (C8) encouraged applicants to consider the risks posed by data biases in AI technologies and required applicants to address data quality, bias mitigation, and ethics in their proposals. Applicants had to consider “risks or issues associated with the proposed timelines and plans for mitigation,” with this criterion accounting for 10% of applicant evaluation.
Some funding calls also described expectations for research novelty more flexibly, ensuring that the exploration of AI-based interventions was suitable for the specific contexts. For example, the Wellcome Mental Health Data Prizes (C8 and C9) clarified that “it is not expected that the team must innovate in their methodology if this is not appropriate for their research.” Similarly, the Climate Change AI Innovation Grants (C13) specified that “it is perfectly acceptable to use a previously published ML algorithm or architecture,” and grounded the exploration of AI on a solid contextual understanding: “Projects should employ or address AI/ML in a way that is well-motivated and well-scoped for the problem setting, which may or may not include the development of novel techniques.”
@@ 4.3. Eligibility Criteria and Team Composition
Funding programs in our corpus commonly viewed interdisciplinary collaborations as valuable for achieving the aims of AI4SG. However, some funding calls had requirements that were biased towards (and could only be realistically met by) teams with existing AI capacities (Section 4.3.1). Funding calls also differed in how they scaffolded involvement of different stakeholders, as well as who was seen as a “domain expert” (e.g., those with scholarly knowledge versus lived experiences). Overall, we observed that many funding documents that encouraged involvement of community stakeholders centered around consultative (rather than inclusive, collaborative, or ownership (Delgado et al., 2023)) modes of participation and only during later AI project stages (i.e., implementation phase) (Section 4.3.1). In Section 4.3.2, we spotlight a few funding documents that went beyond the consultative mode and encouraged community ownership.
@@
Themes along a spectrum of more techno-centric to more balanced approaches related to eligibility and team structures of funding documents in our corpus. More techno-centric: Targeting pre-existing technical capacity while giving less weight to contextual knowledge; Community engagement limited to consultation in later stages of projects. More balanced: Requiring domain expertise or contextual knowledge in project teams; Broadening definitions of “domain experts” and encouraging deeper community engagement beyond consultation.
@@ 4.3.1. More techno-centric
Community engagement limited to consultation in later stages of projects. Across many funding documents in our corpus, we observed that expectations for community engagement—in cases where they existed at all—seldom surpassed the consultation mode, where impacted communities and intended end beneficiaries could provide input, but they were not expected to shape the projects’ overall scope and purpose or play a central role across all stages of the project life cycle. For example, some funding calls only encouraged community consultation during implementation, rather than in the early stages of problem identification and solution development. For example, the Google AI Impact Challenge (C3) specifically asked applicants to “identif[y] the right partners and domain experts needed for implementation.” The funding call explained that involving these stakeholders during implementation could help “deploy the AI model [to support] real-world impact,” where expected community engagement primarily served the purpose of AI deployment.
@@ 4.4. Funders’ Support and Connections
Funding calls in our dataset supported projects ranging from three months to five years (with a median of one and a half years). The longer-term programs were typically funded by governmental agencies (e.g., C17 and C18 by NIH, and C19 by NSF). Overall, we observed a lack of provisions for post-deployment funding support. Furthermore, many funding programs readily provided non-monetary support that was biased towards technical aspects of the projects (Section 4.4.1). Some other funding programs provided broader forms of support beyond the technical, sometimes through partnerships with regional organizations (Section 4.4.2). However, across the funding documents in our corpus, there was a notable lack of guidance and training for applicants to develop skills in community and stakeholder engagement.
@@
Themes along a spectrum of more techno-centric to more balanced approaches related to funders’ support in our corpus. More techno-centric: Lacking post-deployment funding support; Biasing towards technical support. More balanced: Tailoring activities and resources to the needs of applicants beyond technical support; Partnering with regional organizations where the projects are based.
@@ 4.4.2. More balanced
Partnering with regional organizations where the projects are based. A few funding programs involved regional partners in the design of the funding calls (C8-C12). A notable example was the design process of Wellcome Trust’s Mental Health Data Prizes (C8 and C9). After the first round of the data prize in the UK and South Africa (C8), the funder reflected that “it did not attract as many applications as expected from South Africa” (Wellcome Trust, 2023b). After conducting feedback sessions (Wellcome Trust, 2023a), the funder worked with a regional partner, the African Population and Health Research Center, to develop the funding call in Africa (C9). The regional partner played a “primary role” in “prize management” (e.g., developing thematic areas and leading the application selection process), “capacity building” (e.g., providing training in data and mental health research), “matchmaking” (e.g., reach communities and potential participants in Africa), and “landscape analysis of existing mental health research” (e.g., literature review on existing data sources) (C9). The partnership with the regional partner resulted in several key adjustments: providing a literature review on mental health issues in the African context to the applicants, extending the overall timeline for project development, and indicating that a “lived experience expert” would be a member of the application selection panel.
@@ 5. Discussion
Prior literature on “tech for good” establishes that to achieve desired and lasting social impacts, projects typically require a balanced emphasis on both technological capacities and contextual understanding (synthesized in Section 2.2). We characterized such projects as taking a more balanced approach. A bias towards emphasizing technological capacities and potential benefits, with limited contextual knowledge, results in a more techno-centric approach. Using this lens to analyze funding documents, we uncovered a spectrum of conceptual framings of AI4SG and the approaches that funding rhetoric promoted.
@@
Our findings extend prior literature on document analysis of AI imaginaries. We show that some AI4SG funding documents in our corpus similarly reproduce many of the same techno-centric assumptions found in ethics toolkits (Wong et al., 2023b), national AI policies (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022), and military funding solicitations (Widder et al., 2024), but with distinct implications for social-impact work. Like public texts that portray AI as an autonomous and inevitable force (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022; Widder et al., 2024), some AI4SG funding calls in our corpus similarly positioned AI as a transformative solution to large-scale social problems (with limited contextual grounding) and promoted its urgent dissemination to resource-constrained social contexts (like the nonprofit sector and developing countries) (Section 4.1.1). Some funding documents also mirrored AI ethics toolkits that scaffold “technical work for individual technical practitioners” (Wong et al., 2023b), providing little to no guidance on how to involve community stakeholders despite encouraging interdisciplinary collaborations. Yet AI4SG funding documents differ from these domains in their explicit aim to advance social good, making the dissonances between expected outcomes and some techno-centric rhetoric particularly stark. We discuss these dissonances in Section 5.1.
@@
At the same time, a subset of AI4SG documents in our dataset adopted more balanced framings, such as situating AI’s potential within contextual understanding of scoped social problems (Section 4.1.2) and requiring domain expertise or contextual knowledge in core project teams (Section 4.3.2). These examples reveal the presence of funding rhetoric that does encourage community-grounded approaches, a departure from dominant AI imaginaries. While there is complexity in how the content of funding documents influences what actually gets funded, funding rhetoric influences downstream applicants’ behaviors (Peng et al., 2024) and research activities (Widder et al., 2024; Smith et al., 2023). This makes intervening in the design process of funding documents crucial for influencing downstream project approaches before projects are even selected. Building on the more balanced approaches and existing gaps identified in our dataset, we propose opportunities for designing future funding calls for funders, including supporting relationship building, maintenance post-project deployment, and community engagement training for project teams (Section 5.2). Finally, we outline an agenda for meaningful community engagement for funders, and specify the roles that the HCI researchers and practitioners can and should play in helping shape AI4SG funding agendas (Section 5.3)
@@
5.1. Revealing Dissonances Between Intended Outcomes and More Techno-centric Approaches in AI4SG Funding Documents
@@
Many AI4SG funding documents in our dataset expected feasible and sustainable deployment of AI technologies or novel research advancements (Section 4.2). However, the more techno-centric approaches promoted in some funding documents may influence downstream approaches taken by AI4SG project teams, negatively impacting AI4SG’s goal of achieving positive and long-lasting change in real-world contexts (Perera, 2024; Tudor et al., 2024). We highlight three prominent aspects below, in conversation with HCI literature on “tech for good” projects.
@@
First of all, funding documents motivating AI’s use to address large-scale social problems without instantiating concrete contextual benefits (Section 4.1.1) could lead to premature emphasis on scalability, risking putting such a strong emphasis on creating something that works everywhere that it ends up not working particularly well anywhere. As Hanna and Park (2020) note, “scale thinking” necessitates homogenizing users into predefined categories, disregarding the diversity of local contexts. The urgency to scale could incentivize rushed implementations that are poorly suited to diverse contexts (Heeks, 2003, 2002; Dada, 2006). As Burrell and Toyama (2009) put it, “in an effort to solve problems quickly on large scale, a concern for local values, preferences, and opinions may become secondary.” We do not argue that scalable and generalizable approaches should never be pursued, but we suggest that they should be pursued inductively. That is, AI4SG initiatives should focus early efforts on solving specific problems of specific communities effectively and sustainably, and then look for commonalities that could inform the development of generalizable technical capabilities and infrastructures.
@@
Secondly, funding rhetoric that leads with assumptions of AI’s deterministic benefits without acknowledging its risks (Section 4.1.1) could reinforce the wider hype and “enchanted determinism” surrounding AI (Campolo and Crawford, 2020). This narrative fosters unrealistic expectations (Lin et al., 2024; Taylor et al., 2013), promotes techno-solutionism, and diverts attention from underlying social issues (Walsham, 2017; Karusala and Kumar, 2017). Furthermore, when funding documents frame AI as a force for good in the nonprofit sector and in developing countries deterministically (Section 4.1.1), they risk promoting top-down approaches that cast these groups as lacking and in need of catching up (Erete et al., 2023; Valencia, 1997; Dearden and Tucker, 2016). This framing ignores the unique needs (and assets) of the nonprofit sector and developing countries (Hanna and Park, 2020). While corporate donations of computing resources and support (Section 4.4.1) can expand access, they may also reinforce Big Tech’s hegemonic narrative (Collins, 2000; Lukacz, 2024; Nost and Colven, 2022; Seger et al., 2023), addressing surface-level gaps rather than deeper structural issues (Ho et al., 2009b; Burrell and Toyama, 2009; Dearden and Tucker, 2016). Notably, all the funders in our dataset are headquartered in the Global North, even though many funding programs were aimed at applicants or projects located in the Global South. This geographic asymmetry reveals an uneven distribution of hegemonic power (Collins, 2000), as Global North institutions retain disproportionate authority to set funding agendas. These initiatives risk reproducing patterns of saviorism, positioning the Global South as a testing ground for foreign-owned systems while extracting valuable data (Damilare Dosunmu, 2025). These projects also overlook the uneven geography of the AI supply chain; the Global South is already part of AI, but primarily through the underpaid labor that makes AI systems work (Posada, 2022).
@@
5.2. Aligning AI4SG Funding with Intended Outcomes: Recommendations for Funding Call Designs
Building on the more balanced approaches and the existing gaps we identified in our dataset, we offer the following recommendations for funding call designs, placing them in conversation with existing literature on “tech for good” projects. We structure these recommendations around the four key components of funding calls, to serve as a template to inform the design of future funding calls for funders, as well as future HCI research directions in community collaborative approaches.